The International Monetary Fund is trying to raise $900 billion in 6 years for new funds to lend to developing countries. At 6% interest compounded quarterly, how much must it invest today to reach $900 billion in 6 years? (Use the Table provided.)  (Do not round intermediate calculations. Enter your answer in billions of dollar rounded to 2 decimal places.)
Answer #1

nt A = P(1457) 34035 A = 900 t=6 n = 4, re 0.06. 90 = P(1+0.08724 4 P(1.01324 = PX 1.4295 P. 629.59 They hau urut $629.59 Wil